# Elements of Cohomology for model-theorically finite-dimensional groups and Lie algebras [^1] Keywords: nilpotent group, nilpotent Lie algebra, cohomology, finite-dimensional theory MSC 2020 : 17B56, 03C60

### v1

@@ 6 Definable Maschke’s theorem
We will provide another application of the cohomological study of model-theorically finite-dimensional nilpotent groups by proving a definable version of Maschke’s theorem. We will closely follow the argument for the case of ranked groups given by Tindzhogo Ntsiri in [11]. We will need the following definability lemma :
Fact 6.1.
@@
In a finite-dimensional theory, let $V$ be a definable connected $p$-elementary abelian group and let $R$ be an invariant commutative ring of definable endomorphisms of $V$ generated by an infinite definable set $X$. Suppose that $R^{p}=R$. Let $W$ be a definable connected irreducible (in the definable category) sub-$R$-module, such that $V/W$ is also $R$-irreducible (in the definable category). Then $W$ admits a definable $R$-invariant complement.
@@
First, notice that $R$ is unbounded. By considering the action of $R$ on $W$ and $V/W$ and by applying Fact 4.3, the rings $C_{\operatorname{End}_{def}(W)}(R)=K_{W}$ and $C_{\operatorname{End}_{def}(V/W)}(R)=K_{V/W}$ are in fact definable skew-fields. Therefore $\operatorname{ann}_{R}(W)$ and $\operatorname{ann}_{R}(V/W)$ are prime ideals.
We may suppose that $\operatorname{ann}_{R}(W)=\operatorname{ann}_{R}(V/W)$. Indeed, let $0\neq r\in\operatorname{ann}_{R}(W)-\operatorname{ann}(V/W)$. The definable subgroups $\operatorname{im}(r)$ and $W\subseteq\ker(r)$ are $R$-invariant since $R$ is commutative. Then $\ker(r)/W$ and $(\operatorname{im}(r)+W)/W$ are $R$-invariant, and by irreducibility of $V/W$, we have $V=\operatorname{im}(r)+W$. Moreover, $\dim(V)=\dim(\operatorname{im}(r))+\dim(\ker(r))=\dim(\operatorname{im}(r))+\dim(\ker(r))-\dim(W\cap\operatorname{im}(r))$; thus $\dim(W\cap\operatorname{im}(r))=0$ and $\operatorname{im}(r)\cap W$ is trivial by irreducibility of $W$.
Conversely, let $0\neq r\in\operatorname{ann}_{R}(V/W)-\operatorname{ann}_{R}(W)$. Then $\operatorname{im}(r)=W$ and $\ker(r)\cap W$ is trivial by irreducibility of $W$. Therefore $V=W\oplus\ker(r)$.
@@
In $\operatorname{End}_{def}(V)$, we consider the subring $K=R\cdot R^{-1}$; this is a commutative invariant and unbounded field. Let $Z\subseteq K$ be a type-definable subset. For $0\neq v\in V$, the evaluation map $r\mapsto r\cdot v$ is injective since $K$ acts by automorphisms; thus $\dim(Z)=\dim(Z\cdot v)\leq\dim(V)$. By Lemma 6.1, $K$ is in fact a definable field. As a consequence, $W$ is a $K$-subspace of finite linear dimension and admits a definable $R$-invariant complement. ∎
@@
Let $R$ be the subring of $\operatorname{End}_{def}(V)$ generated by $T$. Since $T$ is abelian and $T^{p}=T$, we have $R^{p}=R$. Moreover, $W$ and $V/W$ are $R$-irreducible (in the definable category) : let $W_{1}$ be a finite $R$-sub-module of $W$; in particular, this is a $T$-module, but, by connectedness of $T$, the action is trivial and $W_{1}=0$ since $V^{T}=0$; the proof for $V/W$ is similar since $(V/W)^{T}=0$. Then, we are done by Proposition 6.2. ∎
@@
7 Appendix : Lie algebras
This appendix deals with the case of Lie algebras. Indeed, in [8], Hochshild and Serre defines a cohomology theory for Lie algebras that parallels the theory for groups.
@@ 7 Appendix : Lie algebras
A $\mathfrak{g}$-module is the data of ($\mathfrak{g}$,AA), where $\mathfrak{g}$ is a Lie algebra over a field $K$ and $A$ a $K$-vector space, equipped with a Lie algebra action $\cdot"$ of $\mathfrak{g}$ on $A$. Equivalently, this amounts to giving a Lie algebra homomorphism $\rho$ between $\mathfrak{g}$ and $\operatorname{End}_{K}(A)$.
@@
  • For every integer $n\geq 1$, $C^{n}(\mathfrak{g},A)=\{f\in\operatorname{Hom}(\mathfrak{g}^{n},A):f\text{ is alternating}\}$, where $\operatorname{Hom}(\mathfrak{g}^{n},A)$ denotes the set of applications between $\mathfrak{g}^{n}$ and $A$, which are $K$-linear in each variable (we say also $K$-multilinear). Recall that a $K$-multilinear application $f$ is alternating if $f(x_{1},\dots,x_{n})=0$ whenever two arguments among the $x_{i}$ are equal.
@@
The Lie algebra $\mathfrak{g}$ acts on the cochain groups in a way that is compatible with the differentials :
@@
All the others statements of section 2 are still true when passing to Lie algebras, and the proofs require minor adaptations. In particular, we have the following :
Fact 7.4.
  • Every short exact sequence of $\mathfrak{g}$-modules $0\rightarrow A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ induces a short exact sequence of complexes $0\rightarrow C(\mathfrak{g},A)\overset{i^{*}}{\rightarrow}C(\mathfrak{g},B)\overset{p^{*}}{\rightarrow}C(\mathfrak{g},C)\rightarrow 0$, where $i^{*},p^{*}$ are defined by $i^{*}(f)=i\circ f$ and $p^{*}(f)=p\circ f$.
@@
Remark 7.5.
As pointed out by Moreno Invitti, we can obtain the first six terms of the long exact sequence of cohomology groups induced by a short exact sequence of $\mathfrak{g}$-modules when $\mathfrak{g}$ is an abstract Lie ring. Indeed, it suffices to apply the Snake Lemma to the commutative diagram whose lines are $A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ and $0\rightarrow\operatorname{Der}(\mathfrak{g},A)\overset{i^{*}}{\rightarrow}\operatorname{Der}(\mathfrak{g},B)\overset{p^{*}}{\rightarrow}\operatorname{Der}(\mathfrak{g},C)$. Since we do not use the long exact sequence of cohomology groups for higher values of $n$, all the results of this appendix are still true for Lie rings that are definable in a finite-dimensional theory. Some generalizations of the results of this paper are currently under investigation by Moreno Invitti.
@@
Since $A^{\mathfrak{h}}=A$, by Fact 7.4, the sequence $0\rightarrow H^{1}(\mathfrak{g}/\mathfrak{h},A)\overset{\inf}{\rightarrow}H^{1}(\mathfrak{g},A)\overset{\operatorname{res}}{\rightarrow}H^{1}(\mathfrak{h},A)^{\mathfrak{g}/\mathfrak{h}}$ is exact. Therefore, it suffices to prove that $H^{1}(\mathfrak{h},A)^{\mathfrak{g}/\mathfrak{h}}=0$.
@@
Regarding the adaptation of the argument of section 3 and 4, the only minor variation concerns the abelian case for irreducible module.
@@
Let $\mathfrak{g}$ be a finite-dimensional (in the usual linear sense) Lie algebra and $A$ be a finite-dimensional (in the usual linear sense) $\mathfrak{g}$-module. Suppose that $A$ is $\mathfrak{g}$-irreducible and $A^{\mathfrak{g}}=0$. Then $H^{1}(\mathfrak{g},A)=0$.
@@
We proceed by induction on the length of $A$ in the category of finite-dimensional $\mathfrak{g}$-modules. Since $A$ is finite-dimensional, the length is also finite. If $A=0$, we are done.
First, we way suppose that $\mathfrak{g}$ acts faithfully by Lemma 7.6 and so we have a embedding $\rho:\mathfrak{g}\hookrightarrow\operatorname{End}_{K}(A)$. Consider $R$ the subring of $\operatorname{End}_{K}(A)$ generated by $\mathfrak{g}$ and notice that $A$ is a simple $R$-module. By Schur’s Lemma, $C_{\operatorname{End}_{K}(A)}(R)$ is a skew-field containing $R$; up-to considering its center, we may suppose that it is commutative.
@@
Now, we may generalize the statements of section 4 and section 5 to the case of Lie algebras/Lie rings definable in a finite-dimensional theory. Notice that the Frattini argument obtained for Lie algebras/Lie rings is stronger since the commutativity of the group structure is automatic. In particular, we have the following :
  • In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected nilpotent Lie algebra/Lie ring and $A$ be a definable connected $\mathfrak{g}$-module. Suppose that $A^{\mathfrak{g}}=0$. Then $H^{1}(\mathfrak{g},A)=0$.
  • In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected nilpotent Lie algebra/Lie ring and $A$ be a definable connected $\mathfrak{g}$-module. Suppose that $A^{\mathfrak{g}}=0$. Then for all definable connected sub-modules $V\leq U$ contained in $A$, one has $(U/V)^{\mathfrak{g}}=0$.
  • In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected Lie algebra/Lie ring, and let $\mathfrak{c}$ be a definable connected Cartan subalgebra/subring of a definable connected ideal $\mathfrak{i}$. Then
    $\mathfrak{g}=\mathfrak{i}+N_{\mathfrak{g}}(\mathfrak{c}).$
Some applications of the Frattini argument will be given in an ongoing collaboration with Jules Tindzhogo Ntsiri, that deals with the structure of ranked soluble Lie rings through the analysis of their Cartan subrings (see [12]).
Regarding the adaptation of the content of section 6, we must clarify what semi-simplicity means for Lie algebras in positive characteristic. We will start with the algebra of endomorphisms of a finite-dimensional vector space, for which linear algebra provides a natural notion of semi-simple element. For the following developments, we refer to Chapter II of [6].
@@
The following proposition shows that the two notions of semi-simplicity coincide :
Proposition 7.9.
[6, Proposition 3.3, chap. 2] An element $x$ of $\mathfrak{g}$ is algebraically semi-simple if and only if it induces a semi-simple endomorphism.
@@
In the linear finite-dimensional case, an element of a restricted Lie algebra over a perfect field of characteristic $p>0$ is semi-simple if and only if it belongs to an abelian subalgebra stable under $[p]$, such that $ker([p])=0$ in this subalgebra (see [6, Proposition 3.3 and Theorem 3.7]).
Now, we may use proposition 6.2 to prove the following :
@@
In a finite-dimensional theory, let $\mathfrak{t}$ be a torus and $A$ a definable connected $\mathfrak{t}$-module. Assume that $A$ is abelian $p$-elementary and $\rho(t^{[p]})(x)=\rho(t)^{p}(x)$ for all $t\in\mathfrak{t}$ and $x\in A$, where $\rho$ is the morphism of representation induced by the action of $\mathfrak{t}$. Suppose that $A^{\mathfrak{t}}=0$. Then every definable connected irreducible (in the definable connected category) sub-module of $A$ admits a definable $\mathfrak{t}$-invariant complement.

### v2

@@ 3 Definable Maschke’s theorem
In this section, we provide an application of the cohomological study of finite-dimensional nilpotent groups by proving a definable version of Maschke’s theorem. We will closely follow the argument for the case of ranked groups given by Tindzhogo Ntsiri in [11]. Nevertheless, even for ranked groups, the result presented here is new since we can weaken the hypothesis "decent torus" and replace it by "definable connected abelian p-divisible". We will need the following definability lemma :
Fact 3.1.
@@
In a finite-dimensional theory, let $V$ be a definable connected $p$-elementary abelian group and let $R$ be an unbounded invariant commutative ring of definable endomorphisms of $V$. Suppose that $R^{p}=R$. Let $W$ be a definable connected irreducible sub-$R$-module, such that $V/W$ is also $R$-irreducible. Then $W$ admits a definable $R$-invariant complement.
@@
By considering the action of $R$ on $W$ and $V/W$ and by applying Fact 2.19, the rings $C_{\operatorname{End}_{def}(W)}(R)=K_{W}$ and $C_{\operatorname{End}_{def}(V/W)}(R)=K_{V/W}$ are in fact definable skew-fields. Therefore $\operatorname{ann}_{R}(W)$ and $\operatorname{ann}_{R}(V/W)$ are prime ideals.
We may suppose that $\operatorname{ann}_{R}(W)=\operatorname{ann}_{R}(V/W)$. Indeed, let $0\neq r\in\operatorname{ann}_{R}(W)-\operatorname{ann}(V/W)$. The definable connected subgroups $\operatorname{im}(r)^{\circ}$ and $W\subseteq\ker(r)^{\circ}$ are $R$-invariant since $R$ is commutative. Then $\ker(r)^{\circ}/W$ and $(\operatorname{im}(r)^{\circ}+W)/W$ are $R$-invariant, and by irreducibility of $V/W$, we have $V=\operatorname{im}(r)^{\circ}+W$. Moreover, $\dim(V)=\dim(\operatorname{im}(r)^{\circ})+\dim(\ker(r)^{\circ})=\dim(\operatorname{im}(r)^{\circ})+\dim(\ker(r)^{\circ})-\dim(W\cap\operatorname{im}(r)^{\circ})$; thus $\dim(W\cap\operatorname{im}(r)^{\circ})=0$ and $\operatorname{im}(r)^{\circ}\cap W$ is finite. By linearizing again, $\operatorname{im}(r)^{\circ}\cap W$ is in fact trivial.
Conversely, let $0\neq r\in\operatorname{ann}_{R}(V/W)-\operatorname{ann}_{R}(W)$. Then $\operatorname{im}(r)^{\circ}=W$ by hypothesis and irreducibility of $W$. But $\ker(r)^{\circ}\cap W$ is finite by irreducibility of $W$ and it is in fact trivial by linearizing again. Therefore $V=W\oplus\ker(r)^{\circ}$.
@@
In $\operatorname{End}_{def}(V)$, we consider the subring $K=R\cdot R^{-1}$; this is a commutative invariant and unbounded field. Let $Z\subseteq K$ be a type-definable subset. For $0\neq v\in V$, the evaluation map $K\rightarrow V,k\mapsto r\cdot v$ is injective since $K$ acts by automorphisms; thus $\dim(Z)=\dim(Z\cdot v)\leq\dim(V)$. By Lemma 3.1, $K$ is in fact a definable field. As a consequence, $W$ is a $K$-subspace of finite linear dimension and admits a definable $R$-invariant complement. ∎
@@
Let $R$ be the subring of $\operatorname{End}_{def}(V)$ generated by $T$. Since $T$ is abelian and $T^{p}=T$, we have $R^{p}=R$. Moreover, $W$ and $V/W$ are $R$-irreducible. Then, we are done by Proposition 3.2. ∎
@@
4 The case of Lie rings
4.1 Cohomological results
In [9], Hochshild and Serre defines a cohomology theory for Lie algebras that parallels the theory for groups. The definition can easily be adapted to the more general context of Lie rings; nevertheless, general cohomological results for Lie algebras may fail when passing to Lie rings. Recall that a Lie ring $(\mathfrak{g},+,[,])$ is an abelian group equipped with a bi-additive and anti-symmetric function, $[,]$, that satisfies the Jacobi identity.
@@ 4.1 Cohomological results
A $\mathfrak{g}$-module is the data of ($\mathfrak{g}$,AA), where $\mathfrak{g}$ is a Lie ring and $A$ is an abelian group, equipped with a Lie ring action $\cdot"$ of $\mathfrak{g}$ on $A$. Equivalently, this amounts to giving a Lie ring homomorphism $\rho$ between $\mathfrak{g}$ and $\operatorname{End}(A)$.
@@
  • For every integer $n\geq 1$, $C^{n}(\mathfrak{g},A)=\{f\in\operatorname{Hom}(\mathfrak{g}^{n},A):f\text{ is alternating}\}$, where $\operatorname{Hom}(\mathfrak{g}^{n},A)$ denotes the set of applications between $\mathfrak{g}^{n}$ and $A$, which are $\mathbb{Z}$-linear in each variable (we say also $\mathbb{Z}$-multilinear). Recall that a $\mathbb{Z}$-multilinear application $f$ is alternating if $f(x_{1},\dots,x_{n})=0$ whenever two arguments among the $x_{i}$ are equal.
@@
The Lie ring $\mathfrak{g}$ acts on the cochain groups in a way that is compatible with the differentials :
@@
All the statements that we proved regarding the cohomology of groups are still true when passing to Lie rings, and the proofs require minor adaptations. In particular, we have the following :
Fact 4.4.
  • Every short exact sequence of $\mathfrak{g}$-modules $0\rightarrow A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ induces a "long" exact sequence of cohomology groups of the following form :
    $A^{\mathfrak{g}}\rightarrow B^{\mathfrak{g}}\rightarrow C^{\mathfrak{g}}$
    $\overset{\delta_{0}}{\rightarrow}H^{1}(\mathfrak{g},A)\rightarrow H^{1}(\mathfrak{g},B)\rightarrow\ H^{1}(\mathfrak{g},C)$
@@
Since $A^{\mathfrak{h}}=A$, by Fact 4.4, the sequence $0\rightarrow H^{1}(\mathfrak{g}/\mathfrak{h},A)\overset{\inf}{\rightarrow}H^{1}(\mathfrak{g},A)\overset{\operatorname{res}}{\rightarrow}H^{1}(\mathfrak{h},A)^{\mathfrak{g}/\mathfrak{h}}$ is exact. Therefore, it suffices to prove that $H^{1}(\mathfrak{h},A)^{\mathfrak{g}/\mathfrak{h}}=0$.
@@
There is also a minor variation concerning the abelian case for irreducible module.
@@
In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected Lie ring and $A$ be a definable connected $\mathfrak{g}$-module. Suppose that $A$ is $\mathfrak{g}$-irreducible and $A^{\mathfrak{g}}=0$. Then $H^{1}(\mathfrak{g},A)=0$.
@@
First, we way suppose that $\mathfrak{g}$ acts faithfully by Lemma 4.5 and so we have an embedding $\rho:\mathfrak{g}\hookrightarrow\operatorname{End}_{def}(A)$. Consider $R$ the subring of $\operatorname{End}_{def}(A)$ generated by $\mathfrak{g}$; by Fact 2.19 , $C_{\operatorname{End}_{def}(A)}(R)$ is a definable skew-field containing $R$; up-to considering its center, we may suppose that it is commutative.
@@
Finally, we have the following :
Theorem 4.7.
  • In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected nilpotent Lie ring and $A$ be a definable connected $\mathfrak{g}$-module. Suppose that $A^{\mathfrak{g}}=0$. Then $H^{1}(\mathfrak{g},A)=0$.
  • In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected nilpotent Lie ring and $A$ be a definable connected $\mathfrak{g}$-module. Suppose that $A^{\mathfrak{g}}=0$. Then for all definable connected sub-modules $V\leq U$ contained in $A$, one has $(U/V)^{\mathfrak{g}}=0$.
4.2 Definable Maschke’s theorem for restricted Lie algebras
Regarding Maschke’s theorem, we must pass to Lie algebras since the abstract Lie ring structure seems to be insufficient to capture a good notion of $p$-divisibility/$p$-unipotency. First, we clarify what semi-simplicity means for Lie algebras in positive characteristic. We start with the algebra of endomorphisms of a finite-dimensional vector space, for which linear algebra provides a natural notion of semi-simple element. For the following developments, we refer to Chapter II of [7].
@@ 4.2 Definable Maschke’s theorem for restricted Lie algebras
The following fact shows that the two notions of semi-simplicity coincide :
Fact 4.9.
[7, Proposition 3.3, chap. 2] An element $x$ of $\mathfrak{g}$ is algebraically semi-simple if and only if it induces a semi-simple endomorphism.
@@
In the linear finite-dimensional case, an element of a restricted Lie algebra over a perfect field of characteristic $p>0$ is semi-simple if and only if it belongs to an abelian subalgebra stable under $[p]$, such that $ker([p])=0$ in this subalgebra (see [7, Proposition 3.3 and Theorem 3.7]).
Now, we may use Proposition 3.2 combining with Theorem 4.7 to prove the following :
@@
In a finite-dimensional theory, let $\mathfrak{t}$ be a torus and $V$ a definable connected $\mathfrak{t}$-module. Assume that $V$ is abelian $p$-elementary and $\rho(t^{[p]})(x)=\rho(t)^{p}(x)$ for all $t\in\mathfrak{t}$ and $x\in V$, where $\rho$ is the morphism of representation induced by the action of $\mathfrak{t}$. Suppose that $V^{\mathfrak{t}}=0$. Then $V=V_{1}\oplus\dots\oplus V_{r}$, where each $V_{i}$ is a $\mathfrak{t}$-irreducible definable connected sub-module.
4.3 Frattini’s argument for Lie rings
In the theory of finite groups, $p$-Sylow subgroups play a fundamental role and often allow to obtain structural results about the ambient group. In particular, Frattini’s argument states that, for a finite group $G$, a normal subgroup $H$, and a $p$-Sylow subgroup $S$ of $H$, we have the following :
$G=HN_{G}(S).$
The argument is based on conjugacy results for $p$-Sylow and thus generalizes poorly to the setting of Lie rings : even when one has a suitable analogue of $p$-Sylow subgroups, the "inner" maps are not automorphisms of the Lie ring structure, and passing through the exponential map only partially remedies this defect. However, the conclusion of Frattini’s argument remains valid for finite-dimensional (in the usual linear sense) Lie algebras when $p$-Sylow subgroups are replaced by Cartan subalgebras. The proof relies on the cohomological study of nilpotent Lie algebras (see [3, Theorem 2.1]).
Taking into account our cohomological results and following [3, Theorem 2.1], we are now able to prove an analogous result in the context of finite-dimensional theories.
Definition 4.14.
In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected Lie ring and let $\mathfrak{c}$ be a definable connected nilpotent subring such that
$N_{\mathfrak{g}}(\mathfrak{c})=\mathfrak{c}.$
We then say that $\mathfrak{c}$ is a definable connected Cartan subring.
Proposition 4.15.
In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected Lie ring, and let $\mathfrak{c}$ be a definable connected Cartan subring of a definable connected ideal $\mathfrak{i}$. Then
$\mathfrak{g}=\mathfrak{i}+N_{\mathfrak{g}}(\mathfrak{c}).$
Proof.
We set $\mathfrak{n}=N_{\mathfrak{g}}(\mathfrak{c})$ and we consider the definable connected $\mathfrak{c}$-module $A=\mathfrak{g}/\mathfrak{n}$.
Let us prove that $A^{\mathfrak{c}}=0$ for the adjoint action. Indeed, let $\overline{a}\in A$ be such that $c\cdot\overline{a}=\overline{0}$ for all $c\in\mathfrak{c}$; then $c,a]\in\mathfrak{i}\cap\mathfrak{n}$. But $\mathfrak{i}$ is an ideal and $N_{\mathfrak{i}}(\mathfrak{c})=\mathfrak{c}$, so $c,a]\in\mathfrak{c}$ for all $c$, hence $\overline{a}=\overline{0}$.
Now, consider the $\mathfrak{c}$-module
$(\mathfrak{g}/\mathfrak{n})/((\mathfrak{i}+\mathfrak{n})/\mathfrak{n})\simeq\mathfrak{g}/(\mathfrak{i}+\mathfrak{n}).$
If $(\mathfrak{g}/\mathfrak{n})/((\mathfrak{i}+\mathfrak{n})/\mathfrak{n})\neq 0$, by Theorem 4.7, $\mathfrak{c}$ acts non-trivially, a contradiction since $\mathfrak{c}\leq\mathfrak{i}$ and $\mathfrak{i}$ is an ideal.
Some applications of Frattini’s argument will be given in [12] (in collaboration with Tindzhogo Ntsiri). Notably, we prove that the Frattini subring is nilpotent in the context of ranked connected soluble Lie rings.
A generalization of some parts of this paper is currently under investigation by Moreno Invitti.